For a solution where u = 5.0x10-², calculate Kps from:

a. AgSCN (Kps = 1.1x10-¹²)

b. La (IO3) 3 (Kps = 1.0x10-¹¹)



1
Expert's answer
2021-03-09T06:04:23-0500

Conc = u/1000ml = 5.0×10-2 = 5.0×10-5

a) AgSCN

Mass = 165.95 g ...charge = 1

Kps = Conc 2/(molar mass × charge )

Kps = (25×10-10 )/(165.95×1) = 1.1 ×10-12

b) La(IO3)3

Mass = 663 g ..charge = 3

Kps = (25×10-10) / ( 663×3) = 1.0 × 10-11